# More Credible, Comprehensive and Challenging

### v1

@@ C.2 Difficulty Distribution Analysis
The distribution of difficulty among public datasets and FinanceReasoning is shown in Figure 4. As can be seen, the difficulty distribution of FinanceReasoning surpasses previous datasets with respect to the number of medium and difficult-level questions, thereby facilitating a more accurate reflection of the model’s reasoning ability.
@@ D.1 Example of Financial Function
The constructed financial functions primarily consist of detailed computational code accompanied by comprehensive annotations. These annotations include explanations of function calculations, descriptions of input parameters and output variables, as well as specifications of application scope and constraints. This enhances the readability of the functions and strengthens the applicability of LLMs through well-structured documentation. For detailed information, refer to Table 15.
@@ Appendix G More Experimental Details
The details of the evaluated models are provided in LABEL:tab:model-detail. The token consumption of LLMs and LRMs under CoT and PoT is provided in LABEL:tab:tokens. Additionally, the relationship between token consumption and accuracy across the evaluated models is illustrated in LABEL:fig:cost_effective.
@@
Table 6: Misunderstanding of Problem Error Example by DeepSeek-R1. The wrong answer is highlighted in red, the ground truth is highlighted in green, while the incorrectly generated code is shown in red. The model misunderstood the current 1-year interest rate without risk premium.
Question: What is the approximate FICO score of this individual based on the given factors? Answer to the nearest integer. Question ID: test-2144 Ground Truth: 74 Python Solution:     # Calculate the FICO score     fico_score = (payment_history * payment_history_weight +          accounts_owed * accounts_owed_weight +          length_of_credit_history * length_of_credit_history_weight +          credit_mix * credit_mix_weight +          new_credit * new_credit_weight)     return fico_score Context: A young professional is evaluating his creditworthiness to apply for a mortgage. His financial advisor collects data on his credit behavior and assigns a percentage rating to several aspects of his credit profile: a payment history of 90%, accounts owed at 65%, a length of credit history of 70%, a credit mix at 80%, and new credit at 50%. These factors help approximate a FICO score, which weighs each aspect differently: 35% for payment history, 30% for accounts owed, 15% for the length of credit history, 10% for credit mix, and 10% for new credit. DeepSeek-R1: To calculate the approximate FICO score, each percentage rating is multiplied by its corresponding weight, summed, and then scaled to the FICO range ($300-850$). 1. Payment History (35% weight): $90\%\times 0.35=31.5\%$ 2. Accounts Owed (30% weight): $65\%\times 0.30=19.5\%$ 3. Length of Credit History (15% weight): $70\%\times 0.15=10.5\%$ 4. Credit Mix (10% weight): $80\%\times 0.10=8\%$ 5. New Credit (10% weight): $50\%\times 0.10=5\%$ Total weighted percentage: $31.5\%+19.5\%+10.5\%+8\%+5\%=74.5\%$ Convert to FICO score (scaled to 850): $74.5\%\times 850=0.745\times 850=633.25$ Rounded to the nearest integer: 633. Therefore, the answer is 633.
@@
Table 7: Formula Application Error Example by DeepSeek-R1. The wrong answer is highlighted in red, the ground truth is highlighted in green, while the incorrectly generated code is shown in red. The model lacks knowledge of the FICO score, leading to an incorrect formula.
@@
Question: What is the taxable income from Jane’s ESPP transaction? Answer to the nearest dollar. Question ID: test-2017 Ground Truth: 4200 Python Solution: def solution():  # Define the given variables  purchase_price = 40.0 # Purchase price per share in dollars sale_price = 55.0 # Sale price per share in dollars  shares = 200 # Number of shares  discount = 0.15 # ESPP discount  holding_period = 1.5 # Holding period in years (less than 2)  # Calculate the discounted purchase price per share  discounted_purchase_price = purchase_price * (1 - discount)  # Calculate the total gain from selling the shares  total_gain = (sale_price - discounted_purchase_price) * shares  # Since the holding period is less than 2 years, treat the entire gain as ordinary income  taxable_income = total_gain  # Return the answer rounded to the nearest dollar  return round(taxable_income) # Testing the solution function print(solution()) Context: Jane participates in her company’s Employee Stock Purchase Plan (ESPP), which offers a 15% discount on stock purchases. She bought 200 shares at a purchase price of \$40 per share. After holding the shares for 18 months, she decides to sell them at \$55 per share. The IRS mandates that any share held less than two years must have its gain treated entirely as ordinary income. Jane needs to calculate the taxable income from this transaction to determine her tax liability. DeepSeek-R1: Jane’s taxable income from the ESPP transaction involves two components: the discount received at purchase and the gain from selling the shares, both treated as ordinary income due to the holding period being less than two years. 1. Discount Calculation (Compensation Income): - The purchase price per share (\$40) reflects a 15% discount. The fair market value (FMV) at purchase is calculated as: FMV at $\text{FMV at purchase}=\frac{\$40}{1-0.15}=\frac{\$40}{0.85}\approx\$47.06\text{ per share}.$ Discount per share: \$${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\$47.06}-\$40=\$7.06.$ Therefore, the answer is 3000.
@@
Table 8: Numerical Extraction Error Example by DeepSeek-R1. The wrong answer is highlighted in red, the ground truth is highlighted in green, while the incorrectly generated code is shown in red. The model treated the purchase price (pre-discount price) as the post-discount price.
@@
Question: What is the fair price of the European put option? Answer in USD, rounded to two decimal places. Question ID: test-2208 Ground Truth: 5.46 Python Solution: def solution():  # Define variables  S = 150 # Current stock price in USD  K = 140 # Strike price in USD  r = 0.02 # Risk-free interest rate (as a decimal)  T = 0.5 # Time to maturity in years  volatility = 0.25 # Volatility as a decimal  # Calculate d1 and d2 using the Black-Scholes formula  d1 = (log(S / K) + (r + 0.5 * volatility ** 2) * T) / (volatility * sqrt(T))  d2 = d1 - volatility * sqrt(T)  # Calculate the put option price  put_price = K * exp(-r * T) * norm.cdf(-d2) - S * norm.cdf(-d1)  # Return the put price rounded to two decimal places  return round(put_price, 2) Context: A professional investor is considering purchasing a European put option for a technology stock currently trading at 150 USD. The put option has the following parameters: Current stock price, $S=150$ USD Strike price, $K=140$ USD Time to maturity, $T=0.5$ years Risk-free interest rate, $r=2\%$ per annum Volatility, $\sigma=25\%$ per annum The investor wants to determine the fair price of the put option to make an informed investment decision. DeepSeek-R1: 2. Calculate $d_{1}$ and $d_{2}$: We use the following formulas for $d_{1}$ and $d_{2}$: $d_{1}=\frac{\ln(S/K)+(r+\frac{\sigma^{2}}{2})T}{\sigma\sqrt{T}}=\frac{\ln(150/140)+(0.02+0.25^{2}/2)\cdot 0.5}{0.25\sqrt{0.5}}$ $\ln(150/140)\approx 0.06908$ $(0.02+0.03125)\cdot 0.5=0.025625$ Numerator: $0.06908+0.025625=0.094705$ Denominator: $0.25\cdot\sqrt{0.5}\approx 0.1767767$ $d_{1}\approx\frac{0.094705}{0.1767767}\approx 0.5357$ Next, calculate $d_{2}$: ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}d_{2}=d_{1}-\sigma\sqrt{T}=0.5357-0.1767767\approx 0.3589}$ After refining calculations with precise approximations and standard normal CDF values, the fair price of the European put option rounds to 5.44 USD.
@@
Table 9: Numerical Calculating Error Example by DeepSeek-R1. The wrong answer is highlighted in red, the ground truth is highlighted in green, while the incorrectly generated code is shown in red. The model applied excessive rounding during the calculation steps.
@@
Question: As of december 31 , 2006 , what was the total total cash obligations aggregate carrying value of long-term debt due in 2006 Answer to three decimal places. The Revised Question: As of December 31, 2007, what was the total cash obligations aggregate carrying value of long-term debt due in 2007? Answer to three decimal places. Question ID: test_114 Ground Truth: 9.562 Python Solution: notes_2005 = 338501 total_obligations = 3540009 percent_2005 = notes_2005 / total_obligations answer = percent_2005 * 100 Context:
@@
Table 10: Example of Unsolvable Problem in CodeFinQA. The parts highlighted in red are incorrect, and the parts highlighted in green are the revised ones. Since the question is flawed due to missing 2006 data in the table, it cannot be solved, making the entire output meaningless.
@@
Question: What is the change in total cost of revenue between 2019 and 2018? Answer to two decimal places. The Revised Question: What is the percentage change in total cost of revenue between 2019 and 2018? Answer as a percentage to two decimal places. Question ID: test_142 Ground Truth: -5.81 Python Solution: total_cost_of_revenue_2019 = df["Cost of revenue: – Total cost of revenue"]["2019"] total_cost_of_revenue_2018 = df["Cost of revenue: – Total cost of revenue"]["2018"] answer = ( total_cost_of_revenue_2019 - total_cost_of_revenue_2018 ) / total_cost_of_revenue_2018 * 100.0 Context: Category 2019 2018 Amount Percent Cost of revenue: – Products 29816 34066 -4250 -12% Cost of revenue: – Services 19065 17830 1235 7% Cost of revenue: – Total cost of revenue 48881 51896 -3015 -6%
@@
Table 11: Example of Ambiguous Statement in CodeTAT-QA. The parts highlighted in red are incorrect, and the parts highlighted in green are the revised ones. The ambiguous phrasing of the question leads to multiple possible answers.
@@
Question: Assuming that the variances of the underlying populations are equal, independent samples taken from normally distributed groups display the following features: The combined estimation of the common variance is 2,678.05. What is the suitable t-test statistic to verify the assumption that the average of the two populations are the same? Question ID: validation-24 Ground Truth: 0.938 Topic: Quantitative Analysis & Valuation Python Solution: return (200 - 185) / (2678.05 / 25 + 2678.05 / 18)**0.5 The Revised Python Solution: # Given values sample_size_A = 25 sample_size_B = 18 sample_mean_A = 200 sample_mean_B = 185 combined_variance = 2678.05 # Standard error of the difference in means standard_error = (combined_variance / sample_size_A + combined_variance / sample_size_B)**0.5 # t-test statistic calculation t_statistic = (sample_mean_A - sample_mean_B) / standard_error return t_statistic Context: Sample Size Sample Mean Sample Standard Deviation A 25 200 45 B 18 185 60
@@
Table 12: Example of Oversimplified Process in FinanceMath. The parts highlighted in red are incorrect, and the parts highlighted in green are the revised ones. While the answer is correct, the solution process lacks clarity, and the code omits some steps.
@@
Question: An investor purchases a five-year, 6% annual-coupon payment bond at 89.4535 and sells it in four years. Following the purchase of the bond and prior to the receipt of the first coupon, interest rates go down to 8.5%. What is the realized yield on the bond investment? Answer as a percentage with two decimal places. Before Update Answer: 3.7699999809 Question ID: test-8 Python Solution: coupon_rate = 0.06 initial_price = 89.4535 years_til_maturity = 5 interest_rate = 0.085 coupon = coupon_rate * initial_price discount = initial_price - coupon final_price = discount * (1 + interest_rate) ** (years_til_maturity - 4) realized_yield = (final_price / initial_price) - 1 realized_yield_per = 100.0 * realized_yield round(realized_yield_per, 2) After Update Answer: 8.71 Question ID: test-8 Python Solution: face_value = 100 coupon_rate = 0.06 initial_price = 89.4535 years_til_maturity = 5 interest_rate = 0.085 holding_period = 4 coupon = coupon_rate * face_value sell_price = (face_value + coupon) / ((1 + interest_rate) ** (years_til_maturity - holding_period)) coupon_received = coupon + coupon * (1 + interest_rate) + coupon * (1 + interest_rate) ** 2 + coupon * (1 + interest_rate) ** 3 realized_yield = ((sell_price + coupon_received) / initial_price) ** (1 / holding_period) - 1 realized_yield_per = 100.0 * realized_yield round(realized_yield_per, 2)
@@
Table 13: Example of Incorrect Answer in FinCode. The case had an incorrect answer, which we have corrected. The parts highlighted in red are incorrect, and the parts highlighted in green are the revised ones.
@@
Question: What is the holding period return for the three-year timeframe, given the following annual rates of return for a mutual fund as reported by a researcher (expressed as a percentage)? The Revised Question: What is the holding period return for the three-year timeframe, given the following annual rates of return for a mutual fund as reported by a researcher (expressed as a percentage)? Answer to three decimal places. Question ID: validation-68 Ground Truth: 0.548 Python Solution: return ((1+0.14)*(1-0.10)*(1-0.02)-1)*100 Context: Year Return(%) 2008 14 2009 -10 2010 -2
@@
Financial Function: def calculate_hedge_fund_net_return(initial_investment: float, gross_return_rate: float,  management_fee_rate: float, incentive_fee_rate: float,  hurdle_rate: float) -> float:  """   Calculate the net return for an investor in a hedge fund with fees and hurdle rates.   This function computes the net return in dollar terms for an investor after   accounting for management fees and incentive fees, based on a hurdle rate.   Args:    initial_investment (float): The initial investment amount in millions.    gross_return_rate (float): The gross return rate achieved by the hedge fund.    management_fee_rate (float): The management fee rate as a decimal.    incentive_fee_rate (float): The incentive fee rate as a decimal.    hurdle_rate (float): The hurdle rate as a decimal percentage of initial investment.   Returns:    net_return (float): The net return for the investor in millions to two decimal places.   Notes:    - Applicability: Suitable for hedge funds using a similar fee structure.    - Constraints: Assumes management fee is based on year-end assets and incentive fee     is calculated net of management fee, based on returns over the hurdle rate.    - Considerations: Ensure all rates are provided as decimals (e.g., 2% as 0.02).  """  year_end_assets = initial_investment * (1 + gross_return_rate)  management_fee = year_end_assets * management_fee_rate  net_year_end_assets = year_end_assets - management_fee  hurdle_amount = initial_investment * (1 + hurdle_rate)  excess_return = max(0, net_year_end_assets - hurdle_amount)  incentive_fee = excess_return * incentive_fee_rate  total_fees = management_fee + incentive_fee  net_return = year_end_assets - total_fees - initial_investment  return round(net_return, 2)
Table 15: Example of financial function from function library.

### v2

@@ C.2 Difficulty Distribution Analysis
The distribution of difficulty among public datasets and FinanceReasoning is shown in Figure 4. As can be seen, the difficulty distribution of FinanceReasoning surpasses previous datasets with respect to the number of medium and hard-level questions, thereby facilitating a more accurate reflection of the model’s reasoning ability.
@@ D.1 Example of Financial Function
The constructed financial functions primarily consist of detailed computational code accompanied by comprehensive annotations. These annotations include explanations of function calculations, descriptions of input parameters and output variables, as well as specifications of application scope and constraints. This enhances the readability of the functions and strengthens the applicability of LLMs through well-structured documentation. For detailed information, refer to LABEL:tab:exmFunc.
@@ Appendix G More Experimental Details
The details of the evaluated models are provided in LABEL:tab:model-detail. The token consumption of LLMs and LRMs in CoT and PoT settings is provided in LABEL:tab:tokens. Additionally, the relationship between token consumption and accuracy across the evaluated models is illustrated in LABEL:fig:cost_effective.
@@
Table 6: Misunderstanding of Problem Example by DeepSeek-R1. The wrong answer is highlighted in red, the ground truth is highlighted in green, while the incorrectly generated content is shown in red. At $t=0$, there is no need for risk premium adjustment other than that related to interest rate uncertainty, while the model incorrectly adjusted the discount rate at $t=0$ (by incorporating a risk premium).
Question: Smith demonstrates to Harmonica the total return of a recent transaction performed by the Zenith Fund. Smith explains that the Zenith Fund entered a fully collateralized long position in nearby soybean futures contracts at the quoted futures price of 865.0 (US cents/bushel). Three months afterward, the whole futures position was rolled when the short-term futures price was 877.0 and the long-term futures price was 883.0. What is the Zenith Fund’s three-month total return on the soybean futures transaction, factoring in that the collateral earned an annualized rate of 0.60% during the three-month period between when the initial long position was established and the rolling of the contract (expressed as a percentage)? Answer to three decimal places. Question ID: test-2162 Ground Truth: 0.853 Python Solution: def solution():     …     # Calculating each return     price_return = (current_price_short_term - previous_price) / previous_price     roll_return = (current_price_short_term - current_price_long_term) / current_price_short_term     collateral_return = (months / 12) * (collateral_rate / 100)     # Calculating total return     total_return = price_return + roll_return + collateral_return     return total_return * 100 DeepSeek-R1: The Zenith Fund’s total return consists of the gain from the futures position and the interest earned on the collateral. 1. Futures Return: The futures price increased from 865.0 to 877.0 over three months. $\text{Profit}=877.0-865.0=12.0\ \text{cents/bushel}$ $\text{Return}=\frac{12.0}{865.0}=0.013872\ (\text{or 1.3872\%})$ 2. Collateral Return: The annualized interest rate is 0.60%, so the three-month rate is: $0.60\%\times\frac{3}{12}=0.15\%$ 3. Sum of futures return and collateral return: $1.3872\%+0.15\%=1.5372\%$ Rounded to three decimal places, the total return is 1.537%. Therefore, the answer is 1.537.
@@
Table 7: Formula Application Error Example by DeepSeek-R1. The wrong answer is highlighted in red, the ground truth is highlighted in green, while the incorrectly generated content is shown in red. The model failed to account for the roll yield component in the total return calculation, resulting in an overestimated final answer.
@@
Question: What is the taxable income from Jane’s ESPP transaction? Answer to the nearest dollar. Question ID: test-2017 Ground Truth: 4200 Python Solution: def solution():     # Define the given variables     purchase_price = 40.0 # Purchase price per share in dollars     sale_price = 55.0 # Sale price per share in dollars     shares = 200 # Number of shares     discount = 0.15 # ESPP discount     holding_period = 1.5 # Holding period in years (less than 2)     # Calculate the discounted purchase price per share     discounted_purchase_price = purchase_price * (1 - discount)     # Calculate the total gain from selling the shares     total_gain = (sale_price - discounted_purchase_price) * shares     # Since the holding period is less than 2 years, treat the entire gain as ordinary income     taxable_income = total_gain     # Return the answer rounded to the nearest dollar     return round(taxable_income) Context: Jane participates in her company’s Employee Stock Purchase Plan (ESPP), which offers a 15% discount on stock purchases. She bought 200 shares at a purchase price of \$40 per share. After holding the shares for 18 months, she decides to sell them at \$55 per share. The IRS mandates that any share held less than two years must have its gain treated entirely as ordinary income. Jane needs to calculate the taxable income from this transaction to determine her tax liability. DeepSeek-R1: Jane’s taxable income from the ESPP transaction involves two components: the discount received at purchase and the gain from selling the shares, both treated as ordinary income due to the holding period being less than two years. 1. Discount Calculation (Compensation Income): - The purchase price per share (\$40) reflects a 15% discount. The fair market value (FMV) at purchase is calculated as: $\text{FMV at purchase}=\frac{\$40}{1-0.15}=\frac{\$40}{0.85}\approx\$47.06\text{ per share}.$ - Discount per share: ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\$47.06}-\$40=\$7.06.$ Therefore, the answer is 3000.
@@
Table 8: Numerical Extraction Error Example by DeepSeek-R1. The wrong answer is highlighted in red, the ground truth is highlighted in green, while the incorrectly generated content is shown in red. The model treated the purchase price (pre-discount price) as the post-discount price.
@@
Question: What is the fair price of the European put option? Answer in USD, rounded to two decimal places. Question ID: test-2208 Ground Truth: 5.46 Python Solution: def solution():     …     # Define variables     S = 150 # Current stock price in USD     K = 140 # Strike price in USD     r = 0.02 # Risk-free interest rate (as a decimal)     T = 0.5 # Time to maturity in years     volatility = 0.25 # Volatility as a decimal     # Calculate d1 and d2 using the Black-Scholes formula     d1 = (log(S / K) + (r + 0.5 * volatility ** 2) * T) / (volatility * sqrt(T))     d2 = d1 - volatility * sqrt(T)     # Calculate the put option price     put_price = K * exp(-r * T) * norm.cdf(-d2) - S * norm.cdf(-d1)     # Return the put price rounded to two decimal places     return round(put_price, 2) Context: A professional investor is considering purchasing a European put option for a technology stock currently trading at 150 USD. The put option has the following parameters: Current stock price, $S=150$ USD Strike price, $K=140$ USD Time to maturity, $T=0.5$ years Risk-free interest rate, $r=2\%$ per annum Volatility, $\sigma=25\%$ per annum The investor wants to determine the fair price of the put option to make an informed investment decision. DeepSeek-R1: 2. Calculate $d_{1}$ and $d_{2}$: We use the following formulas for $d_{1}$ and $d_{2}$: $d_{1}=\frac{\ln(S/K)+(r+\frac{\sigma^{2}}{2})T}{\sigma\sqrt{T}}=\frac{\ln(150/140)+(0.02+0.25^{2}/2)\cdot 0.5}{0.25\sqrt{0.5}}$ $\ln(150/140)\approx 0.06908$ $(0.02+0.03125)\cdot 0.5=0.025625$ Numerator: $0.06908+0.025625=0.094705$ Denominator: $0.25\cdot\sqrt{0.5}\approx 0.1767767$ $d_{1}\approx\frac{0.094705}{0.1767767}\approx 0.5357$ Next, calculate $d_{2}$: $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}d_{2}=d_{1}-\sigma\sqrt{T}=0.5357-0.1767767\approx 0.3589$} After refining calculations with precise approximations and standard normal CDF values, the fair price of the European put option rounds to 5.44 USD.
@@
Table 9: Numerical Calculation Error Example by DeepSeek-R1. The wrong answer is highlighted in red, the ground truth is highlighted in green, while the incorrectly generated content is shown in red. The model applied excessive rounding during the calculation steps.
@@
Question: As of december 31 , 2006 , what was the total total cash obligations aggregate carrying value of long-term debt due in 2006 Answer to three decimal places. The Revised Question: As of December 31, 2011, what was the total cash obligations aggregate carrying value of long-term debt due in 2011? Answer as a percentage to three decimal places. Question ID: test-114 Ground Truth: 9.562 Python Solution: def solution():     notes_2011 = 338501     total_obligations = 3540009     percent_2011 = notes_2011 / total_obligations     answer = percent_2011 * 100     return round(answer, 3) Context:
@@
Table 10: Example of Unsolvable Problem in CodeFinQA. The parts highlighted in red are unsolvable , and the parts highlighted in green are the revised ones. Since the question is flawed due to missing 2006 data in the table, it cannot be solved, making the entire output meaningless.
@@
Question: What is the change in total cost of revenue between 2019 and 2018? Answer to two decimal places. The Revised Question: What is the percentage change in total cost of revenue between 2019 and 2018? Answer as a percentage to two decimal places. Question ID: test-142 Ground Truth: -5.81 Python Solution: def solution():     …     total_cost_of_revenue_2019 = df["Cost of revenue: -- Total cost of revenue"]["2019"]     total_cost_of_revenue_2018 = df["Cost of revenue: -- Total cost of revenue"]["2018"]     answer = ( total_cost_of_revenue_2019 - total_cost_of_revenue_2018 ) / total_cost_of_revenue_2018 * 100.0     return round(answer, 2) Context: Category 2019 2018 Amount Percent Cost of revenue: – Products 29816 34066 -4250 -12% Cost of revenue: – Services 19065 17830 1235 7% Cost of revenue: – Total cost of revenue 48881 51896 -3015 -6%
@@
Table 11: Example of Ambiguous Statement in CodeTAT-QA. The parts highlighted in red are ambiguous, and the parts highlighted in green are the revised ones. The ambiguous phrasing of the question leads to multiple possible answers.
@@
Question: Assuming that the variances of the underlying populations are equal, independent samples taken from normally distributed groups display the following features: The combined estimation of the common variance is 2,678.05. What is the suitable t-test statistic to verify the assumption that the average of the two populations are the same? Question ID: validation-24 Ground Truth: 0.938 Topic: Quantitative Analysis & Valuation Python Solution: def solution():     return (200 - 185) / (2678.05 / 25 + 2678.05 / 18)**0.5 The Revised Python Solution: def solution():     # Given values     sample_size_A = 25     sample_size_B = 18     sample_mean_A = 200     sample_mean_B = 185     combined_variance = 2678.05     # Standard error of the difference in means     standard_error = (combined_variance / sample_size_A + combined_variance / sample_size_B)**0.5     # t-test statistic calculation     t_statistic = (sample_mean_A - sample_mean_B) / standard_error     return round(t_statistic, 3) Context: Sample Name Sample Size Sample Mean Sample Standard Deviation A 25 200 45 B 18 185 60
@@
Table 12: Example of Oversimplified Process in FinanceMath. The parts highlighted in red are oversimplified, and the parts highlighted in green are the revised ones. While the answer is correct, the solution process lacks clarity, and the code omits some steps.
@@
Question: An investor purchases a five-year, 6% annual-coupon payment bond at 89.4535 and sells it in four years. Following the purchase of the bond and prior to the receipt of the first coupon, interest rates go down to 8.5%. What is the realized yield on the bond investment? Answer as a percentage with two decimal places. Before Update Question ID: test-8 Answer: 3.7699999809 Python Solution: def solution():     coupon_rate = 0.06     initial_price = 89.4535     years_til_maturity = 5     interest_rate = 0.085     coupon = coupon_rate * initial_price     discount = initial_price - coupon     final_price = discount * (1 + interest_rate) ** (years_til_maturity - 4)     realized_yield = (final_price / initial_price) - 1     realized_yield_per = 100.0 * realized_yield     return round(realized_yield_per, 2) After Update Question ID: test-8 Answer: 8.71 Python Solution: def solution():     face_value = 100     coupon_rate = 0.06     initial_price = 89.4535     years_til_maturity = 5     interest_rate = 0.085     holding_period = 4     coupon = coupon_rate * face_value     sell_price = (face_value + coupon) / ((1 + interest_rate) ** (years_til_maturity - holding_period))     coupon_received = coupon + coupon * (1 + interest_rate) + coupon * (1 + interest_rate) ** 2 + coupon * (1 + interest_rate) ** 3     realized_yield = ((sell_price + coupon_received) / initial_price) ** (1 / holding_period) - 1     realized_yield_per = 100.0 * realized_yield     return round(realized_yield_per, 2)
@@
Table 13: Example of Incorrect Answer in FinCode. The case had an incorrect answer and incorrect python solution, which we have corrected. The parts highlighted in red are incorrect, and the parts highlighted in green are the revised ones.
@@
Question: What is the holding period return for the three-year timeframe, given the following annual rates of return for a mutual fund as reported by a researcher (expressed as a percentage)? The Revised Question: What is the holding period return for the three-year timeframe, given the following annual rates of return for a mutual fund as reported by a researcher (expressed as a percentage)? Answer to three decimal places. Question ID: validation-68 Ground Truth: 0.548 Python Solution: def solution():     return round(((1+0.14)*(1-0.10)*(1-0.02)-1)*100, 3) Context: Year Return (%) 2008 14 2009 -10 2010 -2